OpenStudy (anonymous):

__ Given that QT is an altitude of triangle QRS, M angleQTS=(6x+36), and QR= 2x-5, Find QR.

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

so than QR = 2x-5 = 2*9 -5 = 18-5 = 13 than QR = 13